OpenStudy (anonymous):

Can someone help?!?!?! Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). A. -40x + 70 B. -40x + 64 C. -40x + 24 D. -40x + 31

OpenStudy (anonymous):

@iambatman @across @Hero @TexasBoy123 @texaschic101 @Nurali @Jmmm @ganeshie8 @IMStuck @iGreen

TheSmartOne (thesmartone):

you need to plug in g(x) into f(x)

TheSmartOne (thesmartone):

So -4(10x-6)+7 =f(g(x))

TheSmartOne (thesmartone):

-40x+24+7=f(g(x))

TheSmartOne (thesmartone):

f(g(x))=-40x+31

TheSmartOne (thesmartone):

Which is D.

TheSmartOne (thesmartone):

@AydenCarpenter14

OpenStudy (anonymous):

Thank you @TheSmartOne

TheSmartOne (thesmartone):

You are Welcome! :)